what is the unknown of log1/2 16 + log1/3 9 = b
Answer by stanbon(57219)

what is the unknown of log1/2 16 + log1/3 9 = b
--------------------
log(1/2)(16) = -4
---
log(1/3)9 = -2
-------------------
b = -4-2= -6
======================
Cheers,
Stan H.